OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y= 2/t+5

OpenStudy (anonymous):

solve for t for either the x or y then, plug in the t

OpenStudy (agent0smith):

Solve x = t - 3 for t, then plug that into the y equation

OpenStudy (anonymous):

y=2/x+8?

OpenStudy (agent0smith):

Assuming it's y= 2/(t+5) then yes